OpenStudy (anonymous):

How could you find the limit of x+1/(x+2)(x-3) if the limit goes to -2 , I know that it's going to give us infinity and -infinity but if we take the limit from the right and left it's going to give us - over - and it's +. from the right we take the value -1.9 which going to give us - over + * - " the values".

OpenStudy (kira_yamato):

Use L'hopital's Rule.

OpenStudy (anonymous):

I need you to explain man :( please

OpenStudy (kira_yamato):

So is the expression \[x + \frac{1}{(x+2)(x+3)}?\]

OpenStudy (kira_yamato):

Hi are you there @Cutefriendzoned

OpenStudy (anonymous):

yeah man i'm here , and thank you for helping by the way , but no the expression is like this: lim x goes to -2 (x+1)/(x+2)(x-3)

OpenStudy (kira_yamato):

That makes things easier. Differentiate the numerator and denominator wrt x.

OpenStudy (kira_yamato):

Let L be the limit operator\[L \left( \frac{ u(x) }{ v(x) } \right) = L \left( \frac{ u'(x) }{ v'(x) } \right)\]

OpenStudy (anonymous):

Do we have to differentiate it ? He only wants us to find the limit for left and right. I'm lost now :( what my prof said that just look at (x+2) and see if it gives you + or - if it gives you - and the numerator gives you - then it goes to +infinity. is this correct ?

OpenStudy (kira_yamato):

You differentiate if you wanna use L'hopital's rule. Because in this case you have something indeterminate. So the only way is to apply L'hopital's rule (i.e. differentiate)

OpenStudy (kira_yamato):

\[L \left( \frac{ \color{red}{u(x)} }{ \color{green}{v(x)} } \right) = L \left( \frac{ \color{red}{u'(x)} }{ \color{green}{v'(x)} } \right)\]

OpenStudy (kira_yamato):

Trust me, it will help you quite a bit ^_^

OpenStudy (anonymous):

Oh man , sorry for being late, but it's going to be like this if we're going to do your way lim x goes to 1 which is 1 over lim x goes to -2 for 2x-7 ? either i'm doing something wrong or I didn't get it :(

OpenStudy (tkhunny):

A little late to this party, but do NOT use L'Hospital's Rule on this one. It is NOT an indeterminate form a the rule does not apply. It is the invalid and careless application of L'Hospital's rule. Don't do it! If you mean \( \dfrac{x+1}{(x+2)(x-3)}\), your task is only to determine the sign of anything around x = -2. Clearly, the function increases without bound as it approaches -2. The only question is positive or negative.

OpenStudy (anonymous):

I think that I got it , thank you.

OpenStudy (tkhunny):

* as x approaches -2. I sort of said as the function approaches -2. That would be at least misleading, if not simply incorrect.

OpenStudy (kira_yamato):

Thanks @tkhunny. Guess I overlooked that point on the condition of L'Hopital's

OpenStudy (tkhunny):

No worries. L'Hospital's Rule is a pretty amazing result. It is VERY TEMPTING to use it.

OpenStudy (kira_yamato):

I agree...

OpenStudy (tkhunny):

It is also VERY annoying when the problem statement says, "Find the limit, and DON'T use L'Hospital!" :-) Just make sure it's an indeterminate form before jumping in! Sadly, it doesn't ALWAYS help.

OpenStudy (kira_yamato):

True... Those are the worst problems in limits, ever...